Dublin University Rifle Club

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Adrienne Regan, Head of Sponsorship Sponsorship Management Corporate Relations Eircom 112-114 St.Stephens Green Dublin 2

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Dublin University Rifle Club Sponsorship  1st November 2001

A Chara,

The Dublin University Rifle Club is one of the oldest, largest and most successful sports clubs in the college today. We celebrate the 40^{th} anniversary of the founding of the modern club this year, and last year saw the 160^{th} anniversary of the original college rifle club. We boast 335 members ranging in ages from 18 to 56, making us the fourth largest sporting club in the college.

We currently count the Irish National 50m and 25yd prone rifle champions among our members, DURC members act as part of the National Target Shooting Association committee and we have been requested by the NTSA to manage the logistics for this year’s Air Rifle Nationals.

This past year alone, DURC members have competed in 24 national-level competitions and four international level competitions, including two World Cups; we won a total of 34 prizes at these competitions. DURC has had at least one competitor qualify for the Irish National Squad every year for the past twenty years; at present we have three. This year’s Munich World Cup sees a DURC member acting as Irish Team Manager.

We have an Intervarsity event every year. We have a long history of international competition, and an excellent reputation both here and abroad. A DURC competitor was the first Irish competitor to win a gold medal abroad; we were the first Irish club to win the team trophy in the British Universities Top XX competition, and the first Irish competitor to win an individual medal in that event was from DURC. DURC competitors have represented Ireland at international level many times.

However, our club is seriously underfunded - this summer season for example, the club will be unable to cover the entry fees to competitions for our competitors, who will have to pay their own way. Every year we hold two Open competitions, which are registered with the NTSA, thus allowing competitor’s scores to count towards selection for the national squad. We are unable to afford prizes on an appropriate level for these competitions. Our operational costs are funded to an extent by grants from the Dublin University Central Athletic Club, but there are serious limitations to this funding; Competing outside the country (including Northern Ireland), Prizes, Administrative costs (postage, stationary, phone charges, network access for the club webserver, etc.) are not funded in any way whatsoever. This means that our competition prizes are not of an appropriate standard and our competitive squads must cover all of their own costs when representing the Club (and Ireland) abroad.

We feel that our club is a prime opportunity for corporate sponsorship given our reputation and history. This year being our 40^{th} anniversary, we have several events we hope to run, including a planned formal black-tie dinner function to be held in Trinity College to mark the actual anniversary date. The amount of funding we would require would be on the order of £5,000. This would cover the prizes in competition for the year, including the championship cups for the DURC 10m Air Rifle Open and the DURC 50m Air Rifle Open. Both of these events are Olympic disciplines, and which usually have upwards of forty senior competitors and twenty junior competitors, from all over the country; Target shooting being one of the few sports in Ireland that is organised on a 32-county basis. The competition is also open to competitors from abroad, and this year we are hoping to invite competitors from the Isle of Man and from the UK. We would welcome an “Eircom Cup” in both competitions.

The sponsorship would also allow for a full squad to be fielded by DURC in competitions including the Comber Open, the East Antrim Festival, the British Air Rifle Championships, the Isle of Man Easter Shoot, Bisley Week (the largest smallbore competition in the UK and Ireland), as well as all of the national level competitions.

I would be grateful for the opportunity to discuss this proposal further. I can be contacted at the above addresses; I look forward to hearing from you.

Misé le meas,

Mark Dennehy, DURC Secretary